OpenStudy (anonymous):

evaluate this indefinite integral. Integral ( x^2+1 / x^2) dx

OpenStudy (phi):

is the problem \[ \int \frac{x^2+1}{x^2} \ dx \]?

OpenStudy (anonymous):

yes, It has the brackets around the fraction though

OpenStudy (anonymous):

Divide each term in the numerator by \(x^2\).

OpenStudy (anonymous):

x^2 ?

OpenStudy (anonymous):

ok so x^2/x^2 is 1 right? but 1/x^2?

OpenStudy (anonymous):

Yes. For \(\displaystyle \int\frac{1}{x^2}~dx\), write it as \(x^{-2}\). Now you can simply use the power rule.

OpenStudy (anonymous):

Ok so it would be 1+ x ^ -2?

OpenStudy (anonymous):

and could you please remind me what the power rule is?

OpenStudy (raden):

int x^n dx = 1/(n+1) * x^(n+1) + c that's called the power rule

OpenStudy (anonymous):

Ok I think I got it. thanks